# DynoSurf: Neural Deformation-based Temporally Consistent Dynamic Surface Reconstruction

### v1

@@ Abstract
Figure 1: Temporally-consistent dynamic meshes (i.e., vertices are corresponded and connections are identical over time) reconstructed by our DynoSurf from time-varying 3D point cloud sequences (i.e., no temporal correspondence) without using any shape-prior, ground-truth surface, and ground-truth temporal correspondence. The color and texture map are used to illustrate correspondence across reconstructed mesh frames.
@@ 1 Introduction
With the development of the static surface reconstruction for a single object from a point cloud kazhdan2006poisson ; shen2021deep ; peng2021shape ; huang2023neural , a natural idea to model the temporally consistent dynamic surfaces is to select a keyframe from the point cloud sequences and reconstruct it as a template surface through static reconstruction technology. Then the non-rigid deformation for the template surface is estimated by aligning it with other frames. High requirements are placed on the selection of the keyframe point cloud with fewer missing parts or occlusions and a similar shape to other frames. Some optimization-based dynamic reconstruction frameworks based on fusion continuously update the template shape during frame-by-frame non-rigid tracking newcombe2015dynamicfusion ; dou2016fusion4d ; slavcheva2017killingfusion . These approaches are effective and pragmatic, but most of them use depth as input and cannot handle the sparse point clouds without any camera information. Some network-based method niemeyer2019occupancy ; tang2021learning ; lei2022cadex no longer define or pre-establish the template shape, but jointly learn dynamic reconstruction. This way may achieve better results but increases the difficulty of decoupling the deformation field and the template space, which makes deformation have less practical significance. Moreover, these methods use ground truth to supervise training and have poor generalization. To address these problems, we first design an adaptive keyframe selection strategy and learn the template surface representation, then combine it with a deformation field to form a dynamic reconstruction framework. We reconstruct the temporally consistent surface by joint learning template surface and deformation field.
@@
Furthermore, the design of the deformation field is particularly important. Optimization-based methods li2008global ; li2009robust ; yao2020quasi often employ the embedded deformation graph sumner2007embedded for non-rigid deformation, but they depend on the spatial position and mesh quality of the template surface. Some methods zhang2023self ; park2021nerfies ; tretschk2021non directly use a network to predict the deformation for each point, lacking explicit control of the local structure. bozic2021neural proposes a neural deformation graph to reconstruct dynamic objects. However, they use this graph to reconstruct the geometric surface and non-rigid deformation simultaneously. This makes it need a more complex loss function, leading to the learning unstable. Therefore, we propose a deformation field based on explicit learnable control points, which improves the explicit controllability of the deformation field based on implicit representation. We also learn the blending weights between the points on the template surface and the control points. This allows us to adaptively learn the deformation consistency of surface points based on motion information.
@@
In this paper, we aim to reconstruct a temporally consistent surface from continuous dynamic 3D point cloud sequences. We propose a reconstruction framework combining a template surface representation and a learnable deformation field based on the learnable control points and blending weight. Extensive experiments on three benchmark datasets demonstrate that our proposed temporally-consistent surface reconstruction outperforms the state-of-the-art methods to a large extent.
@@
  • We propose a new learning framework for reconstructing temporally-consistent dynamic surfaces from time-varying point cloud sequences that can be trained without requiring any shape prior, ground-truth surface, and ground-truth temporal correspondence information as supervision.
@@ 2 Related Work
In this section, we briefly review the works closely related to our method, including 3D shape representation, non-rigid shape deformation, and sequential shape reconstruction.
@@
3D Shape Representation.  Generally, 3D shape representation could be divided into two main categories, explicit and implicit, and these can be transformed into one another. The point cloud and triangle mesh are the two most commonly utilized explicit representations for 3D shapes in various applications qi2017pointnet ; li2018pointcnn ; zeng20173dmatch ; hanocka2019meshcnn ; feng2019meshnet . Implicit representation utilizes an isosurface of a field to depict a surface. The Binary Occupancy Field (BOF) kazhdan2013screened ; mescheder2019occupancy ; peng2020convolutional ; boulch2022poco and Signed Distance Field (SDF) kolluri2008provably ; liu2021deep ; park2019deepsdf are two prevalent implicit fields for representing shapes, both dividing the entire space into two regions, inside and outside the shape. The implicit fields could be approximated by a neural network, such as DeepSDF park2019deepsdf , and SAL atzmon2020sal .
@@
Given the shapes explicitly represented as point clouds or triangle meshes, diverse methods exist for their conversion into implicit representations niemeyer2019occupancy ; kazhdan2013screened ; kolluri2008provably ; liu2021deep . Conversely, when transforming implicit shapes into explicit forms, Marching Cubes (MC) MARCHINGCUBE and Marching Tetrahedrons (MT) nielson2008dual stand out as the two widely used methods. Considering the detailed parts of shapes, DMTet shen2021deep and FlexiCubes shen2023flexible introduce offset for each tetrahedron or cube vertex, allowing their methodologies to effectively manage detailed shape components.
@@
Non-rigid Shape Deformation.  Deforming non-rigid shapes presents a challenge due to the high degree of freedom involved in controlling the shape deformation. To mitigate this, methods li2008global ; li2009robust employ an embedded deformation graph sumner2007embedded with local rigid properties for modeling deformation. During optimization, various geometric regularization terms, such as as-conformal-as-possible deformations wu2019global and as-rigid-as-possible deformations yang2019global , are incorporated to impart meaningful deformation. Recently, QNS yao2020quasi and AMM yao2023fast have developed the optimization methods to accelerate the optimization process.
With the advancement of deep learning in 3D vision, numerous learning-based deformation methods have been introduced. DispVoxNet shimada2019dispvoxnets represents source and target shapes using voxel grids, learning an offset field supervised by ground truth offsets. In contrast, FlowNet3D liu2019flownet3d employs PointNet++ qi2017pointnet++ to predict offsets for each point on the shapes. CorrNet3D zeng2021corrnet3d designs a symmetric pipeline to compute deformations between point clouds. RMA-Net feng2021recurrent introduces a new deformation representation with a point-wise combination of several rigid transformations. NDP li2022non proposes a non-rigid motion representation using a pyramid architecture.
@@
Sequential Shape Reconstruction.  Several 3D parametric models, such as SMPL loper2023smpl and SMAL zuffi20173d , have been developed to represent general shapes within specific categories. These parametric models allow for straightforward adjustments in pose and shape by modifying their respective parameters. IPNet chowdhury2022ipnet integrates learned implicit fields with SMPL to control the poses of reconstructed human models. Following this, NPMs palafox2021npms disentangle shapes and poses into latent spaces, enhancing reconstruction accuracy. Conversely, H4D jiang2022h4d and LoRD jiang2022lord initially regress SMPL model parameters to align with each frame of given point clouds. Subsequently, vertex-wise offsets are applied to the resulting SMPL models, approximating the reconstructed shapes to match the provided point cloud sequence. However, these methods are constrained to handling sequences within specific categories, limiting their broad applicability.
@@
Recent studies newcombe2015dynamicfusion ; niemeyer2019occupancy ; tang2021learning ; bozic2021neural have expanded the scope of reconstructing dynamic shapes by predicting deformation fields between different frames. Approaches like Dynamic Fusion newcombe2015dynamicfusion and NDG bozic2021neural establish a deformation graph to capture the deformation fields. O-Flow niemeyer2019occupancy extends ONet mescheder2019occupancy from 3D to 4D, representing deformation fields with a Neural-ODE chen2018neural , while LPDC tang2021learning employs MLPs to model correspondences in parallel. Both O-Flow and LPDC reconstruct the reference shape based on the initial frame, with subsequent frames having their query points mapped to the chosen reference frame. In contrast, I3DMM yenamandra2021i3dmm learns a canonical reference shape for sequences of heads, while CASPR rempe2020caspr and Garment Nets chi2021garmentnets use ground truth canonical coordinates for network training, which is often impractical. Cadex lei2022cadex employs canonical map factorization to represent deformation without requiring correspondences during training, but it may struggle with sequences featuring substantial motions.
@@ 3.1 Problem Statement and Method Overview
Let $\mathcal{P}=\{\mathbf{P}_{1},...,\mathbf{P}_{K}\}$ be a time-varying 3D point cloud sequence capturing the geometric changes or motion of scenes/objects over time, where the correspondence across frames is unknown and $\mathbf{P}_{k}\in\mathbb{R}^{N_{k}\times 3}$ is the $k$-th ($k\in[1,~{}K]$) frame containing $N_{k}$ points. We aim to reconstruct from $\mathcal{P}$ a temporally-consistent mesh sequence =$\mathcal{M}=\{\mathbb{M}_{1},...,\mathbb{M}_{K}\}$, where the $k$-th mesh surface $\mathbb{M}_{k}:=\{\mathbf{V}_{k},\mathcal{E}\}$ with $\mathbf{V}_{k}$ and $\mathcal{E}$ being the vertex set and connection set, respectively, the vertices across meshes are corresponded, and all meshes share the same $\mathcal{E}$.
To address this challenging task, we propose an unsupervised learning framework, named DynoSurf, which requires no ground-truth surface and correspondence information as supervision. As illustrated in Fig. 2, DynoSurf mainly consists of two stages: template surface construction (Sec. 3.2) and deformation-based temporal reconstruction (Sec. 4). Generally, we first select a keyframe from $\mathcal{P}$ and reconstruct its surface based on the deformable tetrahedron representation through a coarse-to-fine learning process. Then, we deform the enhanced template surface to simultaneously align with all frames of $\mathcal{P}$ by learning a control points blending-based deformation field through jointly optimizing the template surface and deformation field. In what follows, we will detail the framework.
Figure 2: Illustration of the proposed DynoSurf, which can reconstruct from time-varying point cloud sequences temporally-consistent dynamic surfaces without requiring any ground-truth surface and temporal correspondence information.
@@ 3.2 Template Surface Representation via Deformable Tetrahedron
In this stage, we aim to construct a template surface, i.e., a mesh fitting the selected point cloud from $\mathcal{P}$, called keyframe. As illustrated in Fig. 3, we initially select a keyframe, denoted as $\mathbf{P}_{k^{*}}\in\mathcal{P}$, and construct its convex hull as a triangle mesh, referred to as $\mathbb{C}$. Subsequently, we process $\mathbb{C}$ to generate a tetrahedron mesh, denoted as $\mathbb{T}$, where the signed distance field (SDF) of the surface underlying $\mathbf{P}_{k^{*}}$ is defined and learned through a coarse-to-fine learning process. The SDF enables us to extract a mesh, denoted as $\overline{\mathbb{M}}:=\{\overline{\mathbf{V}},\mathcal{E}\}$ with $\overline{\mathbf{V}}$ being the vertex set.
@@
Keyframe selection and tetrahedron mesh construction.  Since we will deform the reconstructed template surface from $\mathbf{P}_{k^{*}}$ for temporal reconstruction in the following deformation stage, intuitively, $\mathbf{P}_{k^{*}}$ should be relatively average to facilitate the deformation process. Thus, we determine $k^{*}=\mathop{\arg\min}_{k\in[1,K]}\sum_{l}\texttt{CD}_{\ell_{2}}(\mathbf{P}_{k},\mathbf{P}_{l})$, where $\texttt{CD}_{\ell_{2}}(\cdot,~{}\cdot)$ computes the $\ell_{2}$-norm-based Chamfer distance between two point clouds. We then adopt QuickHull barber1996quickhull to construct the convex hull $\mathbb{C}$ of $\mathbf{P}_{k^{*}}$, which is further dilated and re-meshed to obtain a uniform triangle mesh $\widehat{\mathbb{C}}$ surrounding $\mathbb{C}$. By tetrahedralizing $\widehat{\mathbb{C}}$, we obtain a tetrahedron mesh $\mathbb{T}=\{\mathbf{Q},\mathbf{T}\}$ with $\mathbf{Q}=\{\mathbf{q}_{1},...,\mathbf{q}_{|\mathbf{Q}|}\}$ and $\mathbf{T}=\{\mathbf{T}_{1},...,\mathbf{T}_{|\mathbf{T}|}\}$ being the vertex and tetrahedron sets, respectively. Note that the vertices of $\mathbb{T}$ are distributed both inside and outside of $\mathbb{C}$ due to the dilation and remeshing operations.
Coarse-to-fine learning of deformable tetrahedron-based template surface.  We learn the deformation tetrahedron representation shen2021deep to construct a template surface $\overline{\mathbb{M}}$ that approximates the point surface $\mathbf{P}_{k^{*}}$. Specifically, for each vertex $\mathbf{q}\in\mathbf{Q}$, we train an MLP $F_{\bf{\Phi}}(\cdot)$ parameterized with $\bf{\Phi}$ to learn the SDF value $s(\mathbf{q})$ and the displacement vector $\delta(\mathbf{q})$ with respect to the surface $\overline{\mathbb{M}}$:
@@
where $\Gamma(\cdot)$ is the Position Encoding operator. Then, based on the deformed tetrahedral grid points $\{\mathbf{q}+\delta(\mathbf{q})\}_{\mathbf{q}\in\mathbf{Q}}$ and $\{s(\mathbf{q})\}_{\mathbf{q}\in\mathbf{Q}}$, we can employ the differentiable marching tetrahedron algorithm shen2021deep , denoted as $\texttt{DMT}(\cdot,~{}\cdot)$, to extract $\overline{\mathbb{M}}$, i.e., $\overline{\mathbb{M}}=\texttt{DMT}(\{\mathbf{q}+\delta(\mathbf{q})\},\{s(\mathbf{q})\})$.
To train $F_{\bf{\Phi}}(\cdot)$ to encode the SDF of the surface $\overline{\mathbb{M}}$, we adopt the following coarse-to-fine strategy to make the learning process easier and faster: (1) We initialize $F_{\bf{\Phi}}(\cdot)$ by learning the SDF of $\mathbb{C}$, i.e.,
$\min_{\bf{\Phi}}L_{\textrm{init}1}=\sum_{\mathbf{q}\in\mathbf{Q}}\|s(\mathbf{q})-\texttt{SDF}(\mathbf{q},\mathbb{C})\|^{2},$ (2)
where $s(\mathbf{q})$ is the predicted SDF value of $F_{\bf{\Phi}}(\cdot)$ and $\texttt{SDF}(\mathbf{q},\mathbb{C})$ denotes the ground-truth SDF value of convex hull surface $\mathbb{C}$ at $\mathbf{q}$. (2) We then refine $F_{\bf{\Phi}}(\cdot)$ by constraining $\overline{\mathbb{M}}$ and $\mathbf{P}_{k^{*}}$ to be as close as possible, i.e.,
@@
where $L_{\textrm{CD}}:=\texttt{CD}_{\ell_{1}}(\overline{\mathbb{M}},\mathbf{P}_{k^{*}})$ denotes the $\ell_{1}$-norm-based Chamfer distance between $\overline{\mathbb{M}}$ and $\mathbf{P}_{k^{*}}$; $L_{\textrm{Norm}}^{*}:=\texttt{NC}_{\ell_{1}}(\overline{\mathbb{M}},\mathbf{P}_{k^{*}})$ measures the consistency of the corresponding normals (see the Supplementary Material for the definition of $L_{\textrm{Norm}}^{*}$); $L_{\textrm{SDF}}$ is employed to promote the accuracy of the learned SDF, defined as
@@
where $\texttt{SDF}_{{\rm IMLS}}(\mathbf{q},\mathbf{P}_{k^{*}})$ is the SDF value at $\mathbf{q}$ approximated through implicit moving least-squares kolluri2008provably :
$\texttt{SDF}_{{\rm IMLS}}(\mathbf{q},\mathbf{P}_{k^{*}})=\frac{\sum_{\mathbf{p}_{j}\in\mathcal{N}(\mathbf{q})}\theta(\|\mathbf{q}-\mathbf{p}_{j}\|,{\zeta})\cdot\langle\mathbf{q}-\mathbf{p}_{j},\mathbf{n}_{j}\rangle}{\sum_{\mathbf{p}_{j}\in\mathcal{N}(\mathbf{q})}\theta(\|\mathbf{q}-\mathbf{p}_{j}\|,\zeta)},$ (5)
where $\mathcal{N}(\mathbf{q})$ denotes the set of nearest points in $\mathbf{P}_{k^{*}}$. $\theta(d,\zeta)=\exp(-d^{2}/\zeta^{2})$ is the weight function. We set $|\mathcal{N}(\mathbf{q})|=10$ and $\zeta=0.1$ by default.
Remark.  It is worth noting that the resulting template surface $\overline{\mathbb{M}}$ will be adaptively updated/enhanced during the following deformation stage to facilitate the deformation stage, thus promoting the quality of reconstructed surfaces of the remaining point cloud frames of $\mathcal{P}$.

### v2

@@ Abstract
Keywords:
Dynamic Surface Reconstruction Temporally Consistent Meshes Low-Dimensional Deformation Point Cloud Sequences
Figure 1: Temporally-consistent dynamic meshes (i.e., vertices are corresponded and connections are identical over time) reconstructed by our DynoSurf from continuous dynamic 3D point cloud sequences without using any shape-prior, ground-truth surface, and ground-truth temporal correspondence. The color and texture map are used to illustrate correspondence across reconstructed mesh frames.
@@ 1 Introduction
With the development of the static surface reconstruction for a single object from a point cloud [18, 46, 40, 15], a natural idea to model the temporally consistent dynamic surfaces is to select a keyframe from the point cloud sequences and reconstruct it as a template surface through static reconstruction technology. Then the non-rigid deformation for the template surface is estimated by aligning it with other frames. High requirements are placed on the selection of the keyframe point cloud with fewer missing parts or occlusions and a similar shape to other frames. Some optimization-based dynamic reconstruction frameworks [34, 10, 48] continuously update the template shape during frame-by-frame non-rigid tracking. These approaches are effective and pragmatic, but most of them use depth as input and cannot handle the sparse point clouds without any camera information. Some learning-based method [36, 50, 23] no longer define or pre-establish the template shape, but jointly learn dynamic reconstruction. This way may achieve better results but increases the difficulty of decoupling the deformation field and the template space, which makes deformation have less practical significance. Moreover, these methods use ground truth to supervise training, have poor generalization, and can usually only process sequences of fixed length. To address these problems, we first design an adaptive keyframe selection strategy and learn the template surface representation, then combine it with a deformation field to form a dynamic reconstruction framework.
@@
Furthermore, the design of the deformation field is particularly important. Optimization-based methods [25, 24, 55] often employ the embedded deformation graph [49] for non-rigid deformation, but they depend on the spatial position and mesh quality of the template surface. Some methods [60, 39, 51] directly use a network to predict the deformation for each point, lacking explicit control of the local structure. [6] proposes a neural deformation graph to reconstruct dynamic objects. However, they have some restrictions on the location and connection relationships of nodes, which necessitates the use of multiple loss functions. Therefore, we propose a deformation field based on explicit learnable control points, which improves the explicit controllability of the deformation field based on implicit representation. We also learn the blending weights between the points on the template surface and the control points. This allows us to adaptively learn the deformation consistency of surface points based on motion information.
@@
In this paper, we aim to reconstruct a temporally consistent surface from continuous dynamic 3D point cloud sequences with not-fixed length. We propose a reconstruction framework combining a template surface representation and a learnable deformation field based on the learnable control points and blending weight. Extensive experiments on three benchmark datasets demonstrate that our proposed temporally-consistent surface reconstruction outperforms the state-of-the-art methods to a large extent.
@@
  • We propose a new learning framework for reconstructing temporally-consistent dynamic surfaces from continuous dynamic point cloud sequences that can be trained without requiring any shape prior, ground-truth surface, and ground-truth temporal correspondence information as supervision.
@@ 2 Related Work
We briefly review the most relevant works, including 3D shape representation, non-rigid shape deformation, and sequential shape reconstruction.
@@
3D Shape Representation.  Generally, 3D shape representation could be divided into two main categories, explicit and implicit, and these can be transformed into one another. The point cloud and triangle mesh are the two most commonly utilized explicit representations for 3D shapes in various applications [42, 28, 58, 14, 13]. Implicit representation utilizes an isosurface of a field to depict a surface. The Binary Occupancy Field (BOF) [19, 32, 41, 5] and Signed Distance Field (SDF) [21, 29, 38] are two prevalent implicit fields for representing shapes, both dividing the entire space into two regions, inside and outside the shape. The implicit fields could be approximated by a neural network, such as DeepSDF [38], and SAL [1].
@@
Given the shapes explicitly represented as point clouds or triangle meshes, diverse methods exist for their conversion into implicit representations [36, 19, 21, 29, 22]. Conversely, when transforming implicit shapes into explicit forms, Marching Cubes (MC) [31] and Marching Tetrahedrons (MT) [35] stand out as the two widely used methods. Considering the detailed parts of shapes, DMTet [46] and FlexiCubes [47] introduce offset for each tetrahedron or cube vertex, allowing their methodologies to effectively manage detailed shape components.
@@
Non-rigid Shape Deformation.  Deforming non-rigid shapes presents a challenge due to the high degree of freedom involved in controlling the shape deformation. To mitigate this, [25, 24] employ an embedded deformation graph [49] with local rigid properties for modeling deformation. During optimization, various geometric regularization terms, such as as-conformal-as-possible deformations [53] and as-rigid-as-possible deformations [54], are incorporated to impart meaningful deformation. Recently, [55, 56] have developed optimization methods to accelerate the optimization process. With the advancement of deep learning in 3D vision, numerous learning-based deformation methods have been introduced. CorrNet3D [59] designs a symmetric pipeline to compute deformations between point clouds. RMA-Net [12] introduces a new deformation representation with a point-wise combination of several rigid transformations. NDP [26] proposes a non-rigid motion representation using a pyramid architecture.
@@
Sequential Shape Reconstruction.  Several 3D parametric models, such as SMPL [30] and SMAL [61], have been developed to represent general shapes within specific categories. These parametric models allow for straightforward adjustments in pose and shape by modifying their respective parameters. Subsequently, the combination of neural implicit functions and parametric models enhances the expressive power for reconstruction [37, 17, 16, 7]. However, these methods are constrained to handling sequences within specific categories, limiting their broad applicability.
@@
Recent studies [34, 36, 50, 6] have expanded the scope of reconstructing dynamic shapes by predicting deformation fields between different frames. Approaches like Dynamic Fusion [34] and NDG [6] establish a deformation graph to capture the deformation fields. O-Flow [36] extends ONet [32] from 3D to 4D, representing deformation fields with a Neural-ODE [8], while LPDC [50] employs MLPs to model correspondences in parallel. Both O-Flow and LPDC reconstruct the reference shape based on the initial frame, with subsequent frames having their query points mapped to the chosen reference frame. In contrast, I3DMM [57] learns a canonical reference shape for sequences of heads, while CASPR [44] and Garment Nets [9] use ground truth canonical coordinates for network training, which is often impractical. Cadex [23] employs canonical map factorization to represent deformation without requiring correspondences during training, but it may struggle with sequences featuring substantial motions. Recently, FDN [2] proposes a new computational model that leverages fluid simulation priors to handle dynamic 3D scenes with topological changes.
@@ 3.1 Problem Statement and Method Overview
Let $\{\mathbf{P}_{1},...,\mathbf{P}_{K}\}$ be a continuous dynamic 3D point cloud sequence capturing the geometric changes or motion of scenes/objects over time, where the correspondence across frames is unknown and $\mathbf{P}_{k}\in\mathbb{R}^{N_{k}\times 3}$ is the $k$-th ($k\in[1,~{}K]$) frame containing $N_{k}$ points. We aim to reconstruct from them a temporally-consistent mesh sequence $\{\mathcal{M}_{1},...,\mathcal{M}_{k}\}$ where the $k$-th mesh surface $\mathcal{M}_{k}:=\{\mathbf{V}_{k},\mathbf{F}\}$ with $\mathbf{V}_{k}$ and $\mathbf{F}$ being the vertex set and face set, respectively, the vertices across meshes are corresponded, and all meshes share the same $\mathbf{F}$.
To address this challenging task, we propose an unsupervised learning framework, named DynoSurf, which requires no ground-truth surface and correspondence information as supervision. As illustrated in Fig. 2, DynoSurf mainly consists of two stages: template surface construction (Sec. 3.2) and deformation-based temporal reconstruction (Sec. 4). Generally, we first select a keyframe from the input point clouds, denoted as $\mathbf{P}_{k^{*}}$, and reconstruct its surface based on the deformable tetrahedron representation through a coarse-to-fine learning process. Then, we deform the enhanced template surface to simultaneously align with all point clouds by learning a control points blending-based deformation field through jointly optimizing the template surface and deformation field. In what follows, we will detail the framework.
Figure 2: Illustration of the proposed DynoSurf, which can reconstruct from continuous dynamic point cloud sequences temporally-consistent dynamic surfaces without requiring any ground-truth surface and temporal correspondence information.
@@ 3.2 Template Surface Representation via Deformable Tetrahedron
In this stage, we aim to construct a template surface, i.e., a mesh fitting the selected keyframe point cloud. As illustrated in Fig. 3, we initially select a keyframe $\mathbf{P}_{k^{*}}$, and construct its convex hull as a triangle mesh, referred to as $\mathcal{C}$. Subsequently, we process $\mathcal{C}$ to generate a tetrahedron mesh, denoted as $\mathcal{T}$, where the signed distance field (SDF) of the surface underlying $\mathbf{P}_{k^{*}}$ is defined and learned through a coarse-to-fine learning process. The SDF enables us to extract a mesh, denoted as ¯:=$\overline{\mathcal{M}}:=\{\overline{\mathbf{V}},\mathbf{F}\}$ with $\overline{\mathbf{V}}$ being the vertex set.
@@
Keyframe selection and tetrahedron mesh construction.  Since we will deform the reconstructed template surface from $\mathbf{P}_{k^{*}}$ for temporal reconstruction, intuitively, $\mathbf{P}_{k^{*}}$ should be relatively average to facilitate the deformation process. Thus, we determine $k^{*}=\mathop{\arg\min}_{k\in[1,K]}\sum_{l}\texttt{CD}_{\ell_{2}}(\mathbf{P}_{k},\mathbf{P}_{l})$, where $\texttt{CD}_{\ell_{2}}(\cdot,~{}\cdot)$ computes the $\ell_{2}$-norm-based Chamfer distance between two point clouds.
We then adopt QuickHull [3] to construct the convex hull $\mathcal{C}$ of $\mathbf{P}_{k^{*}}$, which is further dilated and re-meshed to obtain a uniform triangle mesh $\widehat{\mathcal{C}}$ surrounding $\mathcal{C}$. By tetrahedralizing $\widehat{\mathcal{C}}$, we obtain a tetrahedron mesh $\mathcal{T}=\{\mathbf{Q},\mathbf{T}\}$ with $\mathbf{Q}$ and $\mathbf{T}$ being the vertex and tetrahedron sets, respectively. Note that the vertices of $\mathbf{Q}$ are distributed both inside and outside of $\mathcal{C}$ due to the dilation and remeshing operations.
Coarse-to-fine learning of deformable tetrahedron-based template surface.  We learn the deformable tetrahedron representation [46] to construct a template surface $\overline{\mathcal{M}}$ that approximates the point cloud $\mathbf{P}_{k^{*}}$.
@@
where $\Gamma(\cdot)$ is the Position Encoding operator. Then, based on the deformed tetrahedral grid points $\{\mathbf{q}+\delta(\mathbf{q})\}_{\mathbf{q}\in\mathbf{Q}}$ and SDF values $\{s(\mathbf{q})\}_{\mathbf{q}\in\mathbf{Q}}$, we can employ the differentiable marching tetrahedron algorithm [46], denoted as $\texttt{DMT}(\cdot,~{}\cdot)$, to extract $\overline{\mathcal{M}}$, i.e., ¯$\overline{\mathcal{M}}=\texttt{DMT}(\{\mathbf{q}+\delta(\mathbf{q})\},\{s(\mathbf{q})\})$.
To train $F_{\bf{\Phi}}(\cdot)$ to encode the SDF of the surface $\overline{\mathcal{M}}$, we adopt the following coarse-to-fine strategy to make the learning process easier and faster: (1) We initialize $F_{\bf{\Phi}}(\cdot)$ by learning the SDF of $\mathcal{C}$, i.e.,
$\min_{\bf{\Phi}}L_{\textrm{init}1}=\sum_{\mathbf{q}\in\mathbf{Q}}\|s(\mathbf{q})-\texttt{SDF}(\mathbf{q},\mathcal{C})\|^{2},$ (2)
where $s(\mathbf{q})$ is the predicted SDF value of $F_{\bf{\Phi}}(\cdot)$ and $\texttt{SDF}(\mathbf{q},\mathcal{C})$ denotes the ground-truth SDF value of convex hull surface $\mathcal{C}$ at $\mathbf{q}$. (2) We then refine $F_{\bf{\Phi}}(\cdot)$ by constraining $\overline{\mathcal{M}}$ and $\mathbf{P}_{k^{*}}$ to be as close as possible, i.e.,
@@
where $L_{\textrm{CD}}:=\texttt{CD}_{\ell_{1}}(\overline{\mathcal{M}},\mathbf{P}_{k^{*}})$ denotes the $\ell_{1}$-norm-based Chamfer distance between $\overline{\mathcal{M}}$ and $\mathbf{P}_{k^{*}}$; $L_{\textrm{Norm}}^{*}:=\texttt{NC}_{\ell_{1}}(\overline{\mathcal{M}},\mathbf{P}_{k^{*}})$ measures the consistency of the corresponding normals (see the Supplementary Material for details); $L_{\textrm{SDF}}$ is employed to promote the accuracy of the learned SDF, defined as
@@
where $\texttt{SDF}_{{\rm IMLS}}(\mathbf{q},\mathbf{P}_{k^{*}})$ is the SDF value at $\mathbf{q}$ approximated through implicit moving least-squares [21] (refer to the Supplementary Material).
Remark.  It is worth noting that the resulting template surface $\overline{\mathcal{M}}$ will be adaptively updated/enhanced during the following deformation stage to facilitate the deformation stage, thus promoting the quality of reconstructed surfaces of the remaining point cloud frames.